# thanks: corresponding author

### v1

@@ Abstract
We consider an infinite-range interacting quantum spin-1/2 model, undergoing periodic kicking and dissipatively coupled with an environment. In the thermodynamic limit, it is described by classical mean-field equations that can show regular and chaotic regimes. At finite size, we describe the system dynamics using stochastic quantum trajectories. We find that the asymptotic nonstabilizerness (alias the magic, a measure of quantum complexity), averaged over trajectories, mirrors the classical chaotic behavior, while the entanglement entropy has no relation with chaos in the thermodynamic limit.
@@ 1 Introduction
Here we address a rather different scenario, since, in the thermodynamic limit, our model is exactly described by a classical mean-field dynamics, while at finite size it is a quantum many-body spin system. We study the properties of finite-size quantum states along the trajectories and compare them with the corresponding infinite-size classical mean-field dynamics. We focus on the behavior of quantities assessing the many-body quantum coherence, such as the nonstabilizerness and the entanglement entropy.
@@
The nonstabilizerness, also called the magic, has been recently introduced as a quantifier of complexity of a quantum state, complementary to the entanglement [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. It measures the distance from the set of stabilizer states, i. e., “easy” states which can be prepared using a specific class of quantum circuits that are efficiently simulatable with classical algorithms (the Clifford circuits) [29, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49]. We compute it through the stabilizer Rényi entropy [34, 35] averaged over quantum trajectories and using the methods of Ref. [50] for a permutationally invariant system. We find that the magic reaches an asymptotic long-time value displaying a scaling behavior with the system size that is related to the properties of the classical mean-field dynamics. In correspondence of a mean-field chaotic dynamics and in the absence of dissipation, such scaling is logarithmic (i. e., logarithmic with the accessible Hilbert-space dimension). This numerical finding supports a theoretical statement recently put forward in Ref. [51]. Adding dissipation, the scaling becomes power-law, with an exponent showing plateaus for the parameter ranges corresponding to a mean-field chaotic dynamics. The value of the exponent in these plateaus increases with the dissipation strength, while it approaches the zero value (i. e., near to logarithmic scaling) for small dissipation strengths. Thus, for small coupling to the environment, the prediction of Ref. [51] seems to be robust.
@@
On different grounds, the entanglement entropy is known to play a crucial role in the study of many-body dissipative quantum systems evolving along stochastic trajectories. In fact, a flourishing number of recent works has been focusing on local measurements (either discrete or continuous in time) performed in monitored quantum circuits [52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71], as well as in non-interacting [72, 58, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89] and interacting [90, 91, 92, 93, 94, 95, 96, 97] Hamiltonian systems. Moreover, a deep connection between measurement-induced phases and the encoding/decoding properties of a quantum channel has been put forward [98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108]. Situations where the dynamics is only induced by random measurements of non-local string operators have been also considered [109, 110]. In our system we find that, quite unusually, the asymptotic long-time entanglement entropy averaged over trajectories behaves nonmonotonically with the system size: it increases up to a critical size and then starts decreasing. The maximum does not develop any discontinuity or peculiarity at the transition between regular and chaotic behaviors in the classical limit. Thus the entanglement entropy appears to be insensitive to classical chaos in the thermodynamic limit.
@@ 3 Mean-field theory
We notice that, at the mean-field level, the norm $\lVert\vec{m}\rVert=\sqrt{\vec{m}\cdot\vec{m}}$ of $\vec{m}=(m_{x},m_{y},m_{z})$ is conserved. Thus, we can always consider $\lVert\vec{m}\rVert=1$. This allows us to express the magnetization components in terms of just two variables, $Q$ and $P$, such that [119]
@@
where $Q\in[-1,1]$ and $P\in[-\pi/2,\pi/2)$. The inverse relations are $Q=m_{z}$ and $P=\arctan(m_{y}/m_{x})$, except when $Q=1$, where one can consider $P=0$. The variables $(Q,P)$ define the Poincaré section.
@@ 3.1 Phase portraits
Equations (6) and (10) allow us to study the system dynamics at stroboscopic times (i. e., at integer multiples of $\tau$) in a way similar to Poincaré sections [6] [^2] We consider a set of different initial configurations of values $(Q_{0},P_{0})$ that are randomly extracted from uniform distributions covering their respective domains. We consider $N_{\tau}=10^{3}$ periods; in each period, we first evolve the magnetization using Eqs. (6), then apply the kick following Eqs. (10). We store the values $\boldsymbol{(}Q(t_{n}^{+}),P(t_{n}^{+})\boldsymbol{)}$ right after the kick and use them to set the new initial condition for the following period. We repeat this simulation for $M_{0}=300$ initial configurations.. A selection of typical phase-space portraits is shown in Fig. 1, for $h=0.5$ and several choices of $K$ (different columns) and $\gamma$ (different rows). The first column corresponds to $K=0$, that is, to the boundary time crystal model [117]. In this case, when $\gamma=0$, the initial states evolve following closed orbits and the dynamics is always regular. By increasing $\gamma$, we see that one of the fixed points becomes attractive and there is a coexistence of closed orbits and the stable fixed point. When the decay rate increases further [panel (i), for $\gamma=0.5$], closed orbits disappear and there is only one stable fixed point.
@@ 4 Finite-size dynamics along quantum trajectories
First of all, the kicks act at times $t=t_{n}$, by applying to the quantum state the operator $\hat{U}_{K}$ defined in Eq. (3). Between two consecutive kicks, the quantum-jump protocol is the same used in Ref. [97]. We discretize the time in intervals of length $\delta t$. Here the unravelling works by defining the non-Hermitian (nH) Hamiltonian
@@
While averaging over pure-state trajectories is not essential to get expectation values of observables , this procedure becomes crucial whenever evaluating a nonlinear function of the quantum state, such as for the SRE or the EE (see their definitions in the following subsections). For instance, the bipartite von Neumann entropy of the average state $\hat{\rho}$ is not a good entanglement indicator, as it includes classical correlations as well. On the other hand, thanks to the unraveling into pure-state trajectories, it is possible to define the average entanglement of the ensemble as
$\langle\mathcal{E}\rangle=\frac{1}{N_{\text{traj}}}\sum_{k=1}^{N_{\text{traj}}}\mathcal{E}(\ket{\psi_{k}})\neq\mathcal{E}(\hat{\rho}),$ (17)
and similarly goes for the average magic $\langle\mathcal{M}_{2}\rangle$. Numerically one is forced to consider a finite number of trajectories $N_{\text{traj}}$ and this gives rise to an error bar in the average value that is the root mean square deviation divided by $\sqrt{N_{\text{traj}}}$. In the following, when studying dissipative evolutions, we fix $N_{\text{traj}}=1024$, in such a way that the resulting error bars are always of the order of the line width of our plots. As the initial state, we consider the fully polarized eigenstate |$\ket{\psi(0)}=\ket{N/2,N/2}$ of the $\hat{S}_{z}$ total spin component.
Sometimes we are also interested in the asymptotic (long-time) values of such quantities. We estimate them by evolving the system up to long times, until a plateau is reached. In our investigations we find that this saturation is achieved quickly, so that we can stop the dynamics after $N_{\tau}=1000$ periods for the entanglement entropy and $N_{\tau}=100$ for the magic. The value at the plateau always fluctuates, due to the fact that we are considering a finite ensemble of quantum trajectories. Thus, to extract the steady-state behavior, we average over the last $k_{0}$ points, with $k_{0}$ large enough that all transient features have disappeared (specifically, we set $k_{0}=500$ for the EE and $k_{0}=50$ for the SRE). We denote this long-time average by $\overline{\langle\dots\rangle}$ as, for instance,
¯$\overline{\langle\mathcal{E}\rangle}=\frac{1}{N_{\tau}-k_{0}}\sum_{k=k_{0}}^{N_{\tau}}\langle\mathcal{E}(t_{k})\rangle.$ (18)
We also compare dissipative evolutions ($\gamma>0$) with the unitary dynamics case ($\gamma=0$). Unitary results are averaged over $M_{0}=300$ initial random spin coherent states [125]. In fact, the dynamics of a single trajectory entails a source of randomness (the jumps) that is absent in the unitary case. Thus, in the presence of dissipation, the system is able to explore more efficiently the Hilbert space and there is no need for averaging over the different initial conditions.
@@ 4.1 Nonstabilizerness
$\mathcal{M}_{2}(\ket{\psi})=-\ln\Biggl{[}\sum_{\hat{P}\in\mathcal{P}_{N}}\frac{\braket{\psi}{\hat{P}}{\psi}^{4}}{2^{2N}}\Biggr{]}-N\ln 2,$ (19)
@@
[…]
Figure 3 reports the behavior of the nonstabilizerness $\langle\mathcal{M}_{2}(t_{n})\rangle$ versus the stroboscopic time $t_{n}$ in the dissipative quantum kicked top, after averaging over randomness. We first notice that the SRE as a function of time rapidly tends to an asymptotic value. Moreover, when the kick strength $K$ is small and the dynamics is regular, the asymptotic value seems to grow extensively with $N$ (left panels, for $K=0$). On the other hand, for larger kick strengths, the dynamics becomes chaotic and the asymptotic value increases more slowly, possibly as $\ln N$ (right panels, for $K=5$).
To be more quantitative, we first extrapolate the asymptotic value of the SRE as discussed in Sec. 4. We plot it versus the system size (up to $N=256$) in Fig. 4, for various combinations of the kick strength $K$ and the decay rate $\gamma$, corresponding to different values of the Lyapunov exponents of the undarlying semiclassical dynamics [see Fig. 2(a)]. We observe that, for parameters that correspond to regular semiclassical dynamics ($\lambda\leq 0$), the asymptotic SRE is extensive and grows as $\overline{\langle\mathcal{M}_{2}\rangle}\sim N^{\alpha}$, with $\alpha$ close to one. On the other hand, in the chaotic regime, the SRE grows more slowly. Indeed, a numerical fitting returns an exponent that is appreciably smaller than one (see Tab. 1). Given the relatively small available sizes, by simply analyzing these data, we cannot rule out the possibility that the scaling is sub-extensive (but still power law), rather than logarithmic.
Figure 5: The quantity $\delta\mathcal{M}_{N}$ versus the kick strength $K$, for several combinations of $\gamma$, $h$ (see the headers of the panels). The various curves, with different color scale, refer to $N\in\{20,40,80,160\}$, following the same color scheme as in Fig. 3. The background color is related to the value of the largest Lyapunov exponent $\lambda$, following the same color code as in Fig. 2.
To further investigate this point, we define the difference between the asymptotic SRE for a system of size $N$ and that for a system of size $N/2$:
@@
This quantity tends to a constant for increasing $N$, if the SRE increases logarithmically with $N$, while it keeps increasing for a super-logarithmic dependence of the SRE with $N$. In Fig. 5 we plot $\delta\mathcal{M}_{N}$ versus $K$, for different choices of the parameters $\gamma$, $h$ and for various sizes $N$. In the background of each panel, we also show the corresponding value of the largest Lyapunov exponent $\lambda$ in the classical limit, using the same chromatic scale adopted in Fig. 2 (red regions are chaotic, blue regions are regular). In the classically chaotic regions and when $\gamma=0$ [panels (a) and (b)], we clearly see that $\delta\mathcal{M}_{N}$ saturates with $N$. This saturation persists, yet less distinctly, for $\gamma=0.15$ [panels (c), (d)] while it seems to fade away for $\gamma=0.5$ [panels (e), (f)] (although, in this latter case and for the chaotic regions we observe an increase with $N$, this is much slower than in the regular regions). We also see that, whatever the value of $\gamma$, in the classically chaotic regions $\delta\mathcal{M}_{N}$ is almost independent of $K$ (the same occurs for $\overline{\langle\mathcal{M}_{2}\rangle}_{N}$), in contrast with a strong dependence on $K$ that emerges in the classically regular regions. [^5] Results for $\gamma=0$ are more noisy than those for $\gamma>0$ since, as already mentioned, in the unitary case we average over $M_{0}=300$ initial spin-coherent states instead of the $N_{\text{traj}}=1024$ trajectories of the dissipative case.
In Fig. 6, we plot $\delta\mathcal{M}_{N}$ versus the system size $N$, for the same choices of $K$, $\gamma$, and $h$ of Fig. 4. We see that, for the parameters that correspond to a regular mean-field dynamics, $\delta\mathcal{M}_{N}$ strongly depends on $N$, whereas this dependence becomes much weaker when the dynamics is chaotic. We fit these curves with power-law functions $\delta\mathcal{M}_{N}\sim N^{\beta}$ (in this double-logarithmic scale, they appear as straight lines) and the values of $\beta$ resulting from the fit are reported in Tab. 1.
In addition to this, Fig. 7 reports the behavior of $\beta$ versus $K$, for various and fixed values of $\gamma$, $h$. Let us first focus on the classically chaotic regions ($\lambda>0$, red background). We see that, for the unitary case [panels (a) and (b)], $\beta$ vanishes, implying a logarithmic scaling of $\mathcal{M}_{2}$ with $N$ [^6] In the chaotic regime of the unitary case $\gamma=0$, the exponent $\beta$ evaluated numerically only appears (slightly) negative due to noise in the curves shown in Fig. 6.. Since the dimension of the available Hilbert space is $D=N+1$, we have also $\mathcal{M}_{2}\sim\ln D$, thus confirming the prediction of Ref. [51]. In that reference, the focus is on non-symmetric quantum systems made up of $N$ two-level systems, so $D=2^{N}$ and thus $\mathcal{M}_{2}\sim\ln D$ implies an extensive scaling of $\mathcal{M}_{2}$ with $N$.
Figure 6: Same plot as in Fig. 4, but for $\delta\mathcal{M}_{N}$ versus $N$. Dotted lines are the fitting functions $\delta\mathcal{M}_{N}\sim N^{\beta}$.
@@
Table 1: Exponent of the scaling laws $\overline{\langle\mathcal{M}_{2}\rangle}\sim N^{\alpha}$ and $\delta\mathcal{M}_{N}\sim N^{\beta}$ for the parameters shown in Figs. 4 and 6.

### v3

@@ Abstract
We consider an infinite-range interacting quantum spin-1/2 model, undergoing periodic kicking and dissipatively coupled with an environment. In the thermodynamic limit, it is described by classical mean-field equations that can show regular and chaotic regimes. At finite size, we describe the system dynamics using stochastic quantum trajectories. We find that the asymptotic nonstabilizerness (alias the magic, a measure of quantum complexity), averaged over trajectories, mirrors to some extent the classical chaotic behavior, while the entanglement entropy has no relation with chaos in the thermodynamic limit.
@@ 1 Introduction
Here we address a rather different scenario, since, in the thermodynamic limit, our model is exactly described by a classical mean-field dynamics, while at finite size it is a quantum many-body spin system. We study the properties of finite-size quantum states along the trajectories and compare them with the corresponding infinite-size classical mean-field dynamics. We focus on the behavior of quantities assessing the many-body quantum complexity, such as the nonstabilizerness and the entanglement entropy.
@@
The nonstabilizerness, also called the magic, has been recently introduced as a quantifier of complexity of a quantum state, complementary to the entanglement [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42]. It measures the distance from the set of stabilizer states, i. e., “easy” states which can be prepared using a specific class of quantum circuits that are efficiently simulatable with classical algorithms (the Clifford circuits) [29, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52]. We compute it through the stabilizer Rényi entropy [34, 35] averaged over quantum trajectories and using the methods of Ref. [53] for a permutationally invariant system.
We find that the magic reaches an asymptotic long-time value displaying a scaling behavior with the system size that is related to the properties of the classical mean-field dynamics. In correspondence of a mean-field chaotic dynamics and in the absence of dissipation, such scaling is logarithmic (i. e., logarithmic with the accessible Hilbert-space dimension). This numerical finding supports a theoretical statement recently put forward in Ref. [54]. Adding dissipation, the scaling becomes power-law, with an exponent showing plateaus for the parameter ranges corresponding to a mean-field chaotic dynamics. The value of the exponent in these plateaus increases with the dissipation strength, while it approaches the zero value (i. e., near to logarithmic scaling) for small dissipation strengths. Thus, for small decay rate, the prediction of Ref. [54] seems to be robust.
@@
On different grounds, the entanglement entropy is known to play a crucial role in the study of many-body dissipative quantum systems evolving along stochastic trajectories. In fact, a flourishing number of recent works has been focusing on local measurements (either discrete or continuous in time) performed in monitored quantum circuits [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74], as well as in non-interacting [75, 61, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92] and interacting [93, 94, 95, 96, 97, 98, 99, 100, 101] Hamiltonian systems. Moreover, a deep connection between measurement-induced phases and the encoding/decoding properties of a quantum channel has been put forward [102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112]. Situations where the dynamics is only induced by random measurements of non-local string operators have been also considered [113, 114]. In our system we find that, quite unusually, the asymptotic long-time entanglement entropy averaged over trajectories behaves nonmonotonically with the system size: it increases up to a critical size and then starts decreasing. The maximum does not develop any discontinuity or peculiarity at the transition between regular and chaotic behaviors in the classical limit. Thus the entanglement entropy appears to be insensitive to classical chaos in the thermodynamic limit.
@@ 3 Mean-field theory
We notice that, at the mean-field level, the norm $\lVert\vec{m}\rVert=\sqrt{\vec{m}\cdot\vec{m}}$ of $\vec{m}=(m_{x},m_{y},m_{z})$ is conserved. Thus, we can always consider $\lVert\vec{m}\rVert=1$. This allows us to reduce the number of degrees of freedom from three to two and to express the magnetization components in terms of just two variables, $Q$ and $P$, with the change of variables [123]
@@
where $Q\in[-1,1]$ and $P\in[-\pi/2,\pi/2)$. The inverse relations are $Q=m_{z}$ and $P=\arctan(m_{y}/m_{x})$, except when $Q=1$, where one can consider $P=0$. The variables $(Q,P)$ define the Poincaré section and allow for a direct representation of the magnetization dynamics in a two-dimensional space (cf. Fig. 1).
@@ 3.1 Phase portraits
Equations (6) and (10) allow us to study the system dynamics at stroboscopic times (i. e., at integer multiples of $\tau$) in a way similar to Poincaré sections [6] [^2] We consider a set of different initial configurations of values $(Q_{0},P_{0})$ that are randomly extracted from uniform distributions covering their respective domains. We consider $N_{\tau}=10^{3}$ periods; in each period, we first evolve the magnetization using Eqs. (6), then apply the kick following Eqs. (10). We store the values $\boldsymbol{(}Q(t_{n}^{+}),P(t_{n}^{+})\boldsymbol{)}$ right after the kick and use them to set the new initial condition for the following period. We repeat this simulation for $M_{0}=300$ initial configurations.. A selection of typical phase-space portraits is shown in Fig. 1, for $h=0.5$ and several choices of $K$ (different columns) and $\gamma$ (different rows). In all panels, the red dots represent the trajectory of the initial state $\vec{m}(0)=(0,0,1)$, corresponding to $(Q_{0},P_{0})=(1,0)$. The first column corresponds to $K=0$, that is, to the boundary time crystal model [121]. In this case, when $\gamma=0$, the initial states evolve following closed orbits and the dynamics is always regular. By increasing $\gamma$, we see that one of the fixed points becomes attractive and there is a coexistence of closed orbits and the stable fixed point. When the decay rate increases further [panel (i), for $\gamma=0.5$], closed orbits disappear and there is only one stable fixed point.
@@ 4 Finite-size dynamics along quantum trajectories
First of all, the kicks act at times $t=t_{n}$, by applying to the quantum state the operator $\hat{U}_{K}$ defined in Eq. (3). Between two consecutive kicks, the quantum-jump protocol is the same used in Ref. [100]. We discretize the time in intervals of length $\delta t$. Here the unraveling works by defining the non-Hermitian (nH) Hamiltonian
@@
While averaging over pure-state trajectories is not essential to get expectation values of observables, this procedure becomes crucial whenever evaluating a nonlinear function of the quantum state, such as for the SRE or the EE (their definitions are provided in the following subsections). For instance, the bipartite von Neumann entropy of the average state $\hat{\rho}$ is not a good entanglement indicator, as it includes classical correlations as well. In fact, given a generic nonlinear functional $\mathcal{F}$ of $\hat{\rho}$, such as $\mathcal{E}$ or $\mathcal{M}_{2}$, one can define its average value over the ensemble of pure-state trajectories as
$\langle\mathcal{F}\rangle=\frac{1}{N_{\text{traj}}}\sum_{k=1}^{N_{\text{traj}}}\mathcal{F}(\ket{\psi_{k}}\!\bra{\psi_{k}}),$ (17)
which, in general, does not coincide with the evaluation of $\mathcal{F}$ on the state $\hat{\rho}$ itself. That is, contrary to expectations of observables, $\langle\mathcal{F}\rangle\neq\mathcal{F}(\hat{\rho})$.
From a numerical point of view, one is forced to consider a finite number of trajectories $N_{\text{traj}}$. This gives rise to an error bar in the average value that is the root mean square deviation divided by $\sqrt{N_{\text{traj}}}$. In the following, when studying dissipative evolutions, we fix $N_{\text{traj}}=1024$, in such a way that the resulting error bars are always of the order of the line width of our plots. As for the initial state, we consider the fully polarized eigenstate |$\ket{\psi(0)}=\ket{N/2,N/2}$ of the $\hat{S}_{z}$ total spin component.
Sometimes we are also interested in the asymptotic (long-time) values of such quantities. We estimate them by evolving the system up to long times, until a plateau is reached. In our investigations we find that this saturation is achieved quickly, so that we can stop the dynamics after $N_{\tau}=1000$ periods for the entanglement entropy and $N_{\tau}=100$ for the magic. The value at the plateau always fluctuates, due to the fact that we are considering a finite ensemble of quantum trajectories. Thus, to extract the steady-state behavior, we average over the last $k_{0}$ points, with $k_{0}$ large enough that all transient features have disappeared (specifically, we set $k_{0}=500$ for the EE and $k_{0}=50$ for the SRE). We denote this long-time average by $\overline{\langle\dots\rangle}$ such that, given any nonlinear functional $\mathcal{F}$,
¯$\overline{\langle{\mathcal{F}}\rangle}=\frac{1}{N_{\tau}-k_{0}}\sum_{k=k_{0}}^{N_{\tau}}\langle\mathcal{F}(t_{k})\rangle,$ (18)
where $\langle\mathcal{F}(t_{k})\rangle$ is the average of $\mathcal{F}$ over the states of the ensemble at time $t_{k}$. We also compare dissipative evolutions ($\gamma>0$) with the unitary dynamics case ($\gamma=0$). Unitary results are averaged over $M_{0}=300$ initial random spin coherent states [129]. In fact, the dynamics of a single trajectory entails a source of randomness (the jumps) that is absent in the unitary case. Thus, in the presence of dissipation, the system is able to explore more efficiently the Hilbert space and there is no need for averaging over the different initial conditions.
@@ 4.1 Nonstabilizerness
$\mathcal{M}_{2}(\ket{\psi}\!\bra{\psi})=-\ln\Biggl{[}\sum_{\hat{P}\in\mathcal{P}_{N}}\frac{\braket{\psi}{\hat{P}}{\psi}^{4}}{2^{2N}}\Biggr{]}-N\ln 2,$ (19)
@@
Figure 3: The quantity $\delta\mathcal{M}_{N}$ versus the kick strength $K$, for several combinations of $\gamma$, $h$ (see the headers of the panels). The various curves, with different color scale, refer to $N\in\{20,40,80,160\}$, from $N=20$ (darkest) to $N=160$ (lightest). The background color is related to the value of the largest Lyapunov exponent $\lambda$, following the same color code as in Fig. 2.
By looking at the stroboscopic evolution of $\langle\mathcal{M}_{2}\rangle$ averaged over the randomness, we notice that the dynamics converges to plateaus for all $N$ in a short time scale. Interestingly, the magic value at the plateaus seems to scale differently with the system size, depending on the value of $K$. When the kick strength $K$ is small and the dynamics is regular, the asymptotic value seems to grow extensively with $N$. On the other hand, for larger kick strengths, the dynamics becomes chaotic and the asymptotic value increases more slowly, possibly as $\ln N$. These data are shown in Appendix B. To quantitatively clarify this point, here we define the difference between the asymptotic SRE for a system of size $N$ and that for a system of size $N/2$:
@@
This quantity tends to a constant for increasing $N$, if the SRE increases logarithmically with $N$, while it keeps increasing for a super-logarithmic dependence of the SRE with $N$. In Fig. 3 we plot $\delta\mathcal{M}_{N}$ versus $K$, for different choices of the parameters $\gamma$, $h$ and for various sizes $N$. In the background of each panel, we also show the corresponding value of the largest Lyapunov exponent $\lambda$ in the classical limit, using the same chromatic scale adopted in Fig. 2 (red regions are chaotic, blue regions are regular). In the classically chaotic regions and when $\gamma=0$ [panels (a) and (b)], we clearly see that $\delta\mathcal{M}_{N}$ saturates with $N$. This saturation persists, yet less distinctly, for $\gamma=0.15$ [panels (c), (d)] while it seems to fade away for $\gamma=0.5$ [panels (e), (f)] (although, in this latter case and for the chaotic regions we observe an increase with $N$, this is much slower than in the regular regions). We also see that, in the classically chaotic regions, $\delta\mathcal{M}_{N}$ is almost independent of $K$ (the same occurs for $\overline{\langle\mathcal{M}_{2}\rangle}_{N}$, see Appendix B), in contrast with a strong dependence on $K$ that emerges in the classically regular regions. This is especially true for smaller values of $\gamma$, see panels (a-d). [^5] Results for $\gamma=0$ are more noisy than those for $\gamma>0$ since, as already mentioned, in the unitary case we average over $M_{0}=300$ initial spin-coherent states instead of the $N_{\text{traj}}=1024$ trajectories of the dissipative case.
In Fig. 4, we plot $\delta\mathcal{M}_{N}$ versus the system size $N$, for some choices of $K$, $\gamma$, and $h$. We see that, for the parameters that correspond to a regular mean-field dynamics, $\delta\mathcal{M}_{N}$ strongly depends on $N$, whereas this dependence becomes much weaker when the dynamics is chaotic. We fit these curves with power-law functions $\delta\mathcal{M}_{N}\sim N^{\beta}$ (in this double-logarithmic scale, they appear as straight lines) and the values of $\beta$ resulting from the fit are reported in Tab. 1, together with the scaling exponents $\alpha$ of the fitting law $\overline{\langle\mathcal{M}_{2}\rangle}\sim N^{\alpha}$.
Figure 4: The quantity $\delta\mathcal{M}_{N}$ versus $N$. Dotted lines are the fitting functions $\delta\mathcal{M}_{N}\sim N^{\beta}$.
@@
Table 1: Exponent of the scaling laws $\overline{\langle\mathcal{M}_{2}\rangle}\sim N^{\alpha}$ and $\delta\mathcal{M}_{N}\sim N^{\beta}$ for the parameters shown in Fig. 4.
In addition to this, Fig. 5 reports the behavior of $\beta$ versus $K$, for various and fixed values of $\gamma$, $h$. Let us first focus on the classically chaotic regions ($\lambda>0$, red background). We see that, for the unitary case [panels (a) and (b)], $\beta$ vanishes, meaning a logarithmic scaling of $\mathcal{M}_{2}$ with $N$ [^6] In the chaotic regime of the unitary case $\gamma=0$, the exponent $\beta$ evaluated numerically only appears (slightly) negative due to noise in the curves shown in Fig. 4.. Since the dimension of the available Hilbert space is $D=N+1$, this implies $\mathcal{M}_{2}\sim\ln D$. This finding confirms the prediction of Ref. [54]. In that reference, a scaling $\mathcal{M}_{2}\sim\ln D$ is shown to emerge for chaotic quantum systems under unitary dynamics, but the focus is on non-symmetric systems made up of $N$ two-level systems, so $D=2^{N}$, thus implying an extensive scaling of $\mathcal{M}_{2}$ with $N$. In contrast, in our case the dimension of the available Hilbert space is $D=N+1$, thus we have $\mathcal{M}_{2}\sim\ln N$. This observation may make stabilizer-based simulations of such states more feasible, likely due to the unique properties of chaotic states in reduced-dimensional spaces.
At this stage, it is important to stress that permutational invariance does not inherently constrain the maximum achievable SRE, as $\mathcal{M}_{2}^{\text{(max)}}=N\ln 2$ could still be reached even with permutationally invariant states. This bound arises from a flat Pauli spectrum where each Pauli string has an exponentially small weight of $2^{-N}$ over the quantum state—a condition achievable in permutationally invariant states. In fact, in the regular regime, we observe that $\mathcal{M}_{2}$ can scale extensively with $N$ [see, e.g., Fig. 5(e)], demonstrating that this scaling is allowed with permutational invariance as well. That said, our results indicate that, in the chaotic regime, the $N$-scaling of its ensemble average over initial states becomes logarithmic in case of unitary dynamics with a classical chaotic limit [see Fig. 5(a,b)].